OpenStudy (anonymous):

sqrt 2 cos(theta)+1=0

OpenStudy (anonymous):

can you clarify the problem is it:\[\sqrt{2\cos \theta + 1}\]

OpenStudy (anonymous):

\[\sqrt{2} \cos \theta + 1= 0\]

OpenStudy (anonymous):

try to get rid of the 1 that is being added on left side first.. meaning, use the opposite operation of addition.

OpenStudy (anonymous):

after that, get rid of the sqrt 2 that is being multiplied on left side..hence, use the opposite operation of multiplication.

OpenStudy (anonymous):

then, get the inverse cos of both sides, and refer on unit circle for the final answer..